OpenStudy (anonymous):

Which of the following equations could be the result of using the comparison method to solve the system shown? x - 4y - 1 = 0 x + 5y - 4 = 0 4y+ 1 = 4 - 5y -4y+ 1 = 4 - 5y 4y+ 1 = 5y- 4

OpenStudy (anonymous):

-x+4y+1=0 x+5y-4=0 _______ 9y-3=0 y=3/9 ______ x+5(3/9)-4=0 x+15/9-4=0 x-2.3=0 x=2.3 ________ 2.3+5(3/9)-4=0

OpenStudy (anonymous):

4y+ 1 = 5y- 4 is it true